OpenStudy (anonymous):

Solve the equation (4x + 1)(3x - 2) = 0 {-4, 1.5} {-1/4, 2/3} {1/4, 2/3} {-4, 3}

OpenStudy (anonymous):

2nd 1

OpenStudy (lalaly):

x=-1/4 x=2/3 second

OpenStudy (aroub):

-1/4,2/3